The Malayo Colleges Mindanao recently conducted a study analyzing the privacy management habits of teen internet users. In a group of 116 teens, 80 reported having more than 100 followers on Twitter. Find the upper bound of a 90% confidence interval for the true proportion of teens who would report having more than 100 followers on Twitter. 


1
Expert's answer
2022-03-28T15:24:54-0400

"CI=\mu +z\sqrt{\frac{p(1-p)}{n}}"

p=80/116=0.690

"CI=0.69+1.645\sqrt{\frac{0.69(1-0.69)}{116}}=80+0.07=0.76"